# AgenticTagger: Structured Item Representation for Recommendation with LLM Agents

### v1

@@ 7.1. Effect of Multi-agent Self-refinement
Figure 5. Distribution of coverage changes between optimization steps across layers. We report level 2-6, which have enough trials to obtain meaningful results for analysis.
In this ablation, we examine the effect of the iterative refinement loop in AgenticTagger on improving item coverages. To verify its effect, we plot the average percentage differences in item coverage (i.e., items that are assigned at least one feature among the current vocabulary) for each clustering job in Figure 5. As shown, the overall trend is that iterative refinement helps improve item coverage, leading to a reduced number of items that cannot be described by the vocabulary as the number of optimization steps grows. While we report the difference between steps 1 and 2 on Sports dataset, the trend is consist across dataset and optimization steps.
On the other hand, we observe that iterative optimization tends to less steadily bring coverage improvements towards later-layer (i.e., finer-grained) AgenticTagger features. To this end, we hypothesize that to effectively improve item coverage, the LLM must be able to understand increasingly nuanced item-item similarities as the hierarchy depth grows; intuitively, it is easier to separate, for example, high-level differences such as Rock music and Blues music and specific subgenres of Rock music.
7.2. Scaling AgenticTagger to Larger Dataset
In AgenticTagger, a key decision to make when scaling to larger datasets is the balance between controlling the maximum branching factor $b$, i.e., how many features an item can associate with at each granularity level, and how much data to subsample for generating the vocabulary using Algorithm 1. While these two techniques are not mutually exclusive, it is nevertheless helpful to understand which tends to yield better performance for a large dataset.
To investigate this trade-off, we compare the quality of AgenticTagger features generated from a random subset of 20k items with a variant of the same algorithm (i.e., Algorithm 1) where we constrain the maximum branching factor $b$ to be 1. That is, we only consider one randomly selected associated feature at each feature level for an item when generating finer-grained features in the next level. We found that surprisingly, constraining the branching factor yields significantly superior performance compared to subsampling, where the performance drops to a comparable level to P5-CID. We hypothesize that constructing the vocabulary from more items helps AgenticTagger assign valid features to long-tail item groups, which are crucial for capturing niche user interests.
@@ 7.3. Comparison with Free-form Generation in Generative Recommendation
As we discussed earlier, it is hard to directly make use of free-form generated features for generative recommendation, since the method assigns sets of features to items with no constraints on the number of features per item or the “uniqueness” of the features. To illustrate this point, we discuss an earlier attempt that aims at directly using free-form generated features as semantic ID on Sports domain. We choose this domain since it is a larger domain where model performance is known from prior works.
To be specific, we first prompt an LLM to generate free-form descriptors, and treat any ever-generated descriptors as part of the valid vocabulary. We then train a dual-encoder model that predicts which descriptor is generated for each item among the possible descriptor vocabulary (i.e., every descriptor that has been generated by the LLM), following prior work in LLM-based document topic mining (Xie et al., 2025). This process ensures that the potential match between each descriptor-item pair is sufficiently explored, even when the LLM did not generate a particular descriptor initially. We then perform retrieval-augmented generation with another pass of prompting, except that the model this time is asked to select all relevant descriptors from the top 30 suggested descriptors predicted by the aforementioned dual-encoder model. This process results in a maximum of 27, a minimum of 0, and an average of 12.25 descriptors per item.
Due to the large number of descriptors per item, we then prune the descriptors assigned to each item via two methods. (1) Frequency: we split the tags by their occurrence across the corpus into four equal bins, ignoring tags with frequencies smaller than 10 or larger than 2,000, since these are overly rare or common. Each item’s descriptors are selected by taking the most likely tag from each bin, when available, creating a coarse-to-fine tag sequence similar to semantic IDs. (2) K-means: we select 5,000 centroids by running k-means clustering over the text embeddings of the descriptor vocabulary.
However, this framework resulted in poor overall performance: $NDCG@10$ for frequency-based pruning is 0.185, and for k-means-based pruning it is 0.184, which is unlikely to be comparable to modern semantic ID performance. This demonstrates, in contrast, the effectiveness of AgenticTagger in generating hierarchically structured, semantic-ID-like descriptors.
@@ 7.4. Case Study
Item Context Free-form) AgenticTagger
@@
Finally, to understand the improvement brought by AgenticTagger compared to free-form generation, we take an example from our private dataset to analyze the effect of features from both frameworks on connecting the user’s recent interests to the ground-truth item. We analyze a user with two recently interacted items (news feeds, in this case) and the held-out feed content. As shown, the free-form-generated features, while high-quality, failed to generalize across items due to random surface phrase-choices; in contrast, AgenticTagger successfully identified the user’s broad interest in the stock market and detailed interest in stock performance. This performance could be further improved by crossing these features with entity keywords, such as specific stock names and locations.
@@ 8. Conclusion
In this work, we introduce AgenticTagger, a general framework for prompting LLMs to generate natural language item-descriptor representations for recommender systems. We showcase the wide applicability of AgenticTagger across various applications, including term-based retrieval, generative retrieval, ranking, and controllability-oriented tasks. We also discuss the properties of AgenticTagger, such as its complexity and feature cardinality. In the future, we plan to investigate methods for improving effectiveness by introducing AgenticTagger realizations that consider collaborative filtering signals during the feature generation phase, following recent success in automated feature crossing.
@@ A.1. Implementation Details
We directly reuse the data processing implementation open-sourced by Hou et al. 2025b for fair and straightforward comparison. We reuse the reported baseline performance in Hou et al. 2025b, where baseline results for BERTRec, SASRec, FDSA, S³-Rec on Sports and Beauty domain are taken directly from each respective work and rest are re-implementations by Hou et al..
@@
We train the generative recommender model from scratch with AgenticTagger as semantic IDs from scratch for a maximum of 200 epochs, and break the training if there are no improvements on $N@10$ for 20 epochs. Following Hou et al. 2025b, we set training batch size to 256 and tune learning rate from $\{1\times 10^{-3},3\times 10^{-3},5\times 10^{-3}\}$ and weight-decay from $\{0.07,0.1,0.15,0.2\}$. We conduct three repeated experiments with random seeds $\{2025,2026,2027\}$, and report the test set performance with the model checkpoints with best average $N@10$ across runs.
@@
We use standard auto-regressive decoding with a beam size of 20 for inference. However, since AgenticTagger can produce variable length semantic IDs, we do not sample from the base generative recommender for a fixed number of step, and instead generate until we reach the end of sequence token. The generated sequence are then ranked by their beam scores.
@@ A.1.4. AgenticTagger
By default, we choose 15 as the initial cluster size across all datasets and number of refinement cycles to 3. The minimum cluster size for sub-clustering is set to 30. We break the iterative refinement loop if the item coverage for any cluster is greater than 95, or the refiner LLM fail to generate more than 20 tickets. Due to the large number of items for CDs domain, we set the sub-clustering threshold to 100 for branching-controlled version of the algorithm, and 500 for the subsample variant. While the final vocabulary size of AgenticTagger due to stochasticity of program execution, we report the base vocabulary and the number of conflict resolvers (i.e., resolving token in case two items hashed with the same set of AgenticTagger tags) as “base tokens + resolvers” in Table 7 as “vocab_size.” As shown, AgenticTagger is significantly more efficient than its semantic ID counter-parts using less than 3000 words across the board on academia benchmarks.
@@ Appendix B Further Related Works
ML Models with Natural Language Parameters. Beyond recommender systems, there has been continuing effort for introducing parameters tied to explicit natural language concepts into ML models, e.g., concept-bottleneck models (Koh et al., 2020; Sun et al., 2025), typically for its improved interpretability (Zhong et al., 2024). However, this stream of work generally don’t aim for improved accuracy, and frequently assumes there exists a set of concepts (in our case, features) that are readily available (Koh et al., 2020). The most relevant work to ours is perhaps label-free concept bottleneck model (Sun et al., 2025), which prompts a LLM with task description for concepts. However, our aim is to infer a set of features from a large sets items for recommendation, the nature of which is typically hard to infer from the task description itself, unlike domains such as image classification with fixed set of labels.
LLM for Text Mining At Scale Another relevant stream of work to AgenticTagger, as discussed through the paper, is recent advance in leveraging LLMs for scalable text mining, typically minifest as topic/latent theme mining (Xie et al., 2025; Pham et al., 2024) or taxonomy mining (Wan et al., 2024). However, these framework generally do not consider downstream applications, and usually imposes sequential dependencies on the interaction between LLMs and text corpus (Wan et al., 2024; Pham et al., 2024), where as AgenticTagger extensively use parallelize operation, making scaling to recommendation-scale item corpus feasible.
B.1. Effect of Base Models
In our main experiments on public benchmarks, we opt to use the gemini-2.5 family of models since they are released closest to the time the experiments were conducted. However, to investigate if our framework are robust to the choice of based models, we also experimented with gemini-1.5-flash and gemini-1.5-pro as annotator and architect LLM, respectively. As shown in Table 8, both models yields competitive performance, while gemini-2.5 maintains a slight advantage across datasets, likely due to its stronger overall abilities in text generation and understanding.
Table 8. Impact of the backbone model on performance. Gemini-2.5 yields slightly better performance across domains, while Gemini-1.5 yields comparable results.
Backbone LLM Sports Beauty CDs
R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10
B.2. Sample Prompts (Sports Domain)
@@ 3. To ignore outliers:
B.3. Tags Generated
Table 9. Discovered Vocabulary (Sub-sampled) Aross Domains
@@ B.3. Tags Generated
We show qualitative examples of the discovered vocabulary as in Table 9. For presentation purpose, we pick two descriptors per level, but only expand on one of them, until no valid children was found. Note that in AgenticTagger, the depth of each hierarchy path could be variable length, such as we show in the CDs domain example. As showcased in the table, AgenticTagger are able to find clear separations of products, especially towards the earlier layers of the hierarchy (e.g., “Skincare” and “Haircare” products). However, towards the end of the hierarchy, the boundary between products becomes fuzzy, and it is more challenging to come up with clean, disentangled separation rules (e.g., “Specialty Function” and “Ceramic Plate Irons”), and it becomes harder to enforce a nominal constraint on features. While we partially address this issue by instruction the LLM to select best fitting descriptor for each product when multiple are technically correct, developing method to fundamentally address such disentanglement issue remains an open problem. Nevertheless, even at later layers, AgenticTagger could still discover meaning descriptors that separates the products, such as sleeping bags by their applicable temperature range as shown in the Sports domain.
B.4. Effect of Multi-agent Self-refinement on Other Domains
We showcase the improvements of coverage by reporting the distribution of delta changes in coverage rate (i.e., the number of items that can be described by at least one descriptor) for each clustering job $Q$ in Figure 6. The trend is consistent across domains, where (1) our designed multi-agent self-refinement loop effectively refines the vocabulary, improving item coverage; and (2) the trend of improvement (i.e., delta greater than zero) becomes less stable in later layer, possibility due to it is more challenging to generate disentangled descriptors, as we qualitatively analyzed in Section B.3. Regardless, wilcoxon signed-rank test show improvements are significant (i.e., delta greater than zero) with p-value smaller than 0.05, except for level five and six on CDs domain, corroborating the general effectiveness of the self-refinement mechanism.

### v2

@@ 6.1. Effect of Multi-agent Self-refinement
Figure 4. Distribution of coverage changes between optimization steps across layers. We report level 2 to level 6, which have enough trials to obtain meaningful results for analysis.
In this ablation, we examine the effect of the iterative refinement loop in AgenticTagger on improving item coverage. To verify its effect, we plot the average percentage differences in item coverage (i.e., items that are assigned at least one feature among the current vocabulary) for each clustering job in Figure 4. As shown, the overall trend is that iterative refinement helps improve item coverage, leading to a reduced number of items that cannot be described by the vocabulary as the number of optimization steps grows. While we report the difference between steps 1 and 2 on the Sports dataset, the trend remains consistent across datasets and optimization steps.
On the other hand, we observe that iterative optimization tends to bring less steady coverage improvements to later-layer (i.e., finer-grained) AgenticTagger features. To this end, we hypothesize that to effectively improve item coverage, the LLM must be able to understand increasingly nuanced item-to-item similarities as the hierarchy depth increases; intuitively, it is easier to separate high-level differences, such as Rock music and Blues music, than specific subgenres of Rock music.
6.2. Scaling AgenticTagger to Larger Datasets
In AgenticTagger, a key decision to make when scaling to larger datasets is the balance between controlling the maximum branching factor $b$, i.e., how many features an item can be associated with at each granularity level, and how much data to subsample for generating the vocabulary using Algorithm 1. While these two techniques are not mutually exclusive, it is nevertheless helpful to understand which tends to yield better performance for a large dataset.
To investigate this trade-off, we compare the quality of AgenticTagger features generated from a random subset of 20k items with that of the original scaling choice we made on the CDs and Vinyl domain, where we constrained the branching factor to 1. That is, we only consider one randomly selected associated feature at each feature level for an item when generating finer-grained features in the next level. We found that, surprisingly, constraining the branching factor yields significantly superior performance compared to subsampling, where the performance drops to a level comparable to P5-CID. We hypothesize that constructing the vocabulary from more items helps AgenticTagger assign valid features to long-tail item groups, which are crucial for capturing niche user interests.
@@ 6.3. Comparison with Free-form Generation in Generative Recommendation
As discussed earlier, it is difficult to directly use free-form generated features for generative recommendation, as this method assigns sets of features to items with no constraints on vocabulary choice. To illustrate this point, we discuss an earlier attempt aimed at directly using free-form generated features as semantic IDs in the Sports domain.
We first prompt an LLM to generate free-form descriptors and treat all ever-generated descriptors as part of the valid vocabulary. We then train a dual-encoder model that predicts which descriptors are generated for each item among the possible descriptor vocabulary (i.e., every descriptor that has been generated by the LLM), following prior work in LLM-based document topic mining (Xie et al., 2025), where positive signals come from LLM generations. This process ensures that the potential match between each descriptor-item pair is sufficiently explored, even when the LLM did not generate a particular descriptor initially. Finally, we perform retrieval-augmented descriptor generation with another pass of prompting, except that the LLM this time is asked to select all relevant descriptors from the top 30 suggested descriptors predicted by the aforementioned dual-encoder model. This process yields a maximum of 27, a minimum of 0, and an average of 12.25 descriptors per item.
Due to the large number of descriptors per item, we then prune the descriptors assigned to each item via two methods. (1) Frequency: we split the tags by their occurrence across the corpus into four equal bins, ignoring tags with frequencies smaller than 10 or larger than 2,000, since these are overly rare or common. Each item’s descriptors are then selected by taking the most frequent tag from each popularity bin, where available, creating a coarse-to-fine tag sequence similar to semantic IDs. (2) K-means: we select 5,000 centroids by running K-means clustering over the text embeddings of the descriptor vocabulary, and swap the descriptors with their assigned centroid IDs, reducing the number of descriptors per item.
However, this framework results in poor overall performance: $\text{NDCG}@10$ for frequency-based pruning is 0.185, and for K-means-based pruning, it is 0.184, neither of which is comparable to modern semantic ID performance. This demonstrates, in contrast, the effectiveness of AgenticTagger in generating hierarchically structured, semantic-ID-like descriptors.
Table 4. Comparison on AgenticTagger performance with different LLMs on generative recommendation.
Backbone Sports Beauty CDs
N@5 N@10 N@5 N@10 N@5 N@10
Gemini-1.5 0.0184 0.0237 0.0319 0.0397 0.0349 0.0439
Gemini-2.5 0.0194 0.0251 0.0332 0.0408 0.0358 0.0453
6.4. Performance with Different LLMs
In our generative recommendation experiments on public benchmarks, we opted to use the gemini-2.5 family of models, as they were released closest to the time the experiments were conducted. However, to investigate if our framework is robust to the choice of base models, we also experimented with gemini-1.5-flash and gemini-1.5-pro as annotator and architect LLMs, respectively. In our experiments, we find both models yield competitive performance, while gemini-2.5 maintains a slight advantage across datasets, likely due to its stronger overall abilities in text generation and understanding.
@@ 6.5. Case Study
Item Context Free-form AgenticTagger
@@
Finally, to understand the improvement brought by AgenticTagger compared to free-form generation, we took an example from our private dataset to analyze the effect of features from both frameworks on connecting the user’s recent interests to the ground-truth item. As shown, the free-form-generated features, while of generally high coherence, failed to generalize across items due to random surface phrase choices; in contrast, AgenticTagger successfully identified the user’s broad interest in the stock market and detailed interest in stock performance. While in both cases LLMs are effective at extracting general user interest in stocks, AgenticTagger is the only method that can organically connect user history to the ground-truth interacted item, demonstrating the value of vocabulary-constrained, structured descriptor generation.
@@ 7. Conclusion
In this work, we introduce AgenticTagger, a general framework for prompting LLMs to generate natural language item-descriptor representations for recommender systems. We showcase the wide applicability of AgenticTagger across various applications, including generative retrieval, ranking, and controllability-oriented, critique-based recommendation. We also discuss the properties of AgenticTagger, such as its complexity and feature cardinality. In the future, we plan to investigate methods for improving effectiveness by introducing AgenticTagger implementations that consider collaborative filtering signals during the feature generation phase, following the recent success of automated feature crossing.
@@ A.1. Implementation Details
We directly reuse the data processing implementation open-sourced by Hou et al. 2025b for fair and straightforward comparison. We reuse the reported baseline performance in Hou et al. 2025b, where baseline results for BERTRec, SASRec, FDSA, and S³-Rec on the Sports and Beauty domains are taken directly from each respective work, and the rest are re-implementations by Hou et al..
@@
We train the generative recommender model from scratch with AgenticTagger as semantic IDs for a maximum of 200 epochs, breaking the training if there are no improvements in $N@10$ for 20 epochs. Following Hou et al. 2025b, we set training batch size to 256 and tune learning rate from $\{1\times 10^{-3},3\times 10^{-3},5\times 10^{-3}\}$ and weight-decay from $\{0.07,0.1,0.15,0.2\}$. We conduct three repeated experiments with random seeds $\{2025,2026,2027\}$, and report the test set performance with the model checkpoints with best average $N@10$ across runs.
@@
We use standard auto-regressive decoding with a beam size of 20 for inference. However, since AgenticTagger can produce variable length semantic IDs, we do not sample from the base generative recommender for a fixed number of steps; instead, we generate until we reach the end-of-sequence token. The generated sequences are then ranked by their beam scores.
@@ A.1.4. AgenticTagger
By default, we set the initial cluster size to 15 across all datasets and the number of refinement cycles to 3 (while the architect LLM can choose to exceed this number). The minimum cluster size for sub-clustering is set to 30, i.e., any group of more than 30 items would trigger the procedure for producing sub-category descriptors for the said group. We break the iterative refinement loop if the item coverage for any cluster is greater than 95 or if the refiner LLM fails to generate more than 20 error cases. Due to the large number of items in the CDs domain, we set the sub-clustering threshold to 100 for the branching-controlled version of the algorithm and 500 for the subsampled variant. While the final vocabulary size of AgenticTagger varies due to the stochasticity of program execution, we report the base vocabulary and the number of conflict resolvers (i.e., resolving tokens used when two items hash to the same set of AgenticTagger tags) as “base tokens + resolvers” in Table 7 under “vocab_size.” As shown, AgenticTagger is significantly more efficient than its semantic ID counterparts, using fewer than 3,000 words across the board on academic benchmarks.
@@ Appendix B Further Related Works
ML Models with Natural Language Parameters. Beyond recommender systems, there have been continuing efforts to introduce parameters tied to explicit natural language concepts into ML models, e.g., concept-bottleneck models (Koh et al., 2020; Sun et al., 2025), typically for their improved interpretability (Zhong et al., 2024). However, this stream of work generally does not aim for improved accuracy and frequently assumes there exist sets of concepts (in our case, features) that are readily available (Koh et al., 2020). The most relevant work to ours is perhaps the label-free concept bottleneck model (Sun et al., 2025), which prompts an LLM with a task description for concepts. However, our aim is to infer a set of features from large sets of items for recommendation, the nature of which is typically difficult to infer from the task description itself, unlike domains such as image classification with a fixed set of labels.
LLM for Text Mining At Scale Another relevant stream of work to AgenticTagger, as discussed throughout the paper, consists of recent advances in leveraging LLMs for scalable text mining, typically manifesting as topic/latent theme mining (Xie et al., 2025; Pham et al., 2024) or taxonomy mining (Wan et al., 2024). However, these frameworks generally do not consider downstream applications and usually impose sequential dependencies on the interaction between LLMs and text corpora (Wan et al., 2024; Pham et al., 2024), whereas AgenticTagger uses parallelized operations, making scaling to recommendation-scale item corpora feasible.
LLM-generated Semantic IDs Concurrent to this work, we note there are also recent explorations of LLM-generated natural language semantic IDs, dubbed Term IDs (Zhang et al., 2026), where one prompts an LLM for free-form generated semantic IDs over a batch of similar items for adapting pre-trained LLMs for recommendation. To this end, Term IDs impose fewer constraints on the vocabulary itself, since a pre-trained LLM is able to generalize information across semantically similar Term IDs. On the other hand, AgenticTagger aims to develop a descriptor vocabulary with minimal constraints on the downstream model. To illustrate this point, AgenticTagger enables a 2-layer transformer model (the same as Tiger (Rajput et al., 2023)) to perform better than the 0.4B variant of LLM-based GR with Term IDs on the Sports and Beauty domains, showcasing the effectiveness of AgenticTagger.
B.1. Sample Prompts (Sports Domain)
@@ 3. To ignore outliers:
B.2. Ranking Performance on Smaller Dataset
Metric Item ID (Baseline) AgenticTagger (Ours)
R@5 0.0833 $\pm$ 0.0068 0.1205 $\pm$ 0.0080
N@5 0.0589 $\pm$ 0.0051 0.0853 $\pm$ 0.0060
R@10 0.1163 $\pm$ 0.0079 0.1751 $\pm$ 0.0093
N@10 0.0695 $\pm$ 0.0052 0.1028 $\pm$ 0.0061
In this section, we discuss further performance comparisons in ranking experiments, where we report both Recall and NDCG in Table 8. This subset contains 1,791 users, 15,830 items, and 47,260 interactions, making the collaborative filtering signal less salient due to the smaller dataset size. As shown, our method consistently outperforms the ID-based method by a larger margin than in the main experiment, verifying our intuition: when the dataset is large and the collaborative filtering signal is more stable, assigning more item-specific parameters helps in memorizing item-specific information, sometimes leading to better ranking performance. Conversely, when the dataset is sparse, content-aware structured discrete features shared among items help convergence significantly, making AgenticTagger features potentially applicable to cold-start settings, where they lead to better generalization.
B.3. Cardinality of Free-form Generation
To investigate whether free-form generation indeed leads to vocabulary explosion, we conduct an ablation study where we prompt an LLM to generate three keywords over the private dataset discussed in Section 5. Indeed, there were approximately 90k unique keywords generated over 210k documents; however, only 21k (or 24%) of the descriptors appeared in more than one item. Consequently, these features are inapplicable for natively acting as compact, ID-like features, making downstream modeling challenging. On the other hand, descriptors generated by AgenticTagger have a much higher utilization rate (87%) across a vocabulary of over 2k, demonstrating their effectiveness.
B.4. Tags Generated
We show qualitative examples of the discovered vocabulary in Table 9. For presentation purposes, we pick two descriptors per level but only expand on one of them until no valid children are found. Note that in AgenticTagger, the depth of each hierarchy path can be of variable length, as shown in the CDs domain example. As showcased in the table, AgenticTagger is able to find clear separations of products, especially towards the earlier layers of the hierarchy (e.g., “Skincare” and “Haircare” products). However, towards the end of the hierarchy, the boundary between products becomes fuzzy, and it is more challenging to come up with clean, disentangled separation rules (e.g., “Specialty Function” and “Ceramic Plate Irons”), making it harder to enforce a nominal constraint on features. While we partially address this issue by instructing the LLM to select the best-fitting descriptor for each product when multiple are technically correct, developing methods to fundamentally address such disentanglement issues remains an open problem. Nevertheless, even at later layers, AgenticTagger can still discover meaningful descriptors that separate the products, such as sleeping bags by their applicable temperature range, as shown in the Sports domain.
B.5. Effect of Multi-agent Self-refinement on Other Domains
We showcase the improvements in coverage by reporting the distribution of delta changes in the coverage rate in Figure 5. The trend is consistent across domains: (1) our designed multi-agent self-refinement loop effectively refines the vocabulary, improving item coverage; and (2) the trend of improvement (i.e., delta greater than zero) becomes less stable in later layers, possibly because it is more challenging to generate disentangled descriptors, as qualitatively analyzed in Section B.4. Regardless, the Wilcoxon signed-rank test shows that improvements are significant (i.e., delta greater than zero) with p-values smaller than 0.05, except for levels five and six in the CDs domain, corroborating the general effectiveness of the self-refinement mechanism.
Table 9. Discovered Vocabulary (Sub-sampled) Across Domains
@@ B.5. Effect of Multi-agent Self-refinement on Other Domains
Table 10. Summary of Notations
Notation Description
$\mathcal{U},\mathcal{I}$ Sets of users and items
$S_{u}$ Interaction history for user $u\in\mathcal{U}$
$X_{i}$ Raw attributes for item $i\in\mathcal{I}$
$\Pi,\pi$ LLMs, and a specific instance
$\mathcal{E}$ Set of natural language error feedback
AA Descriptor feature vocabulary
$a$ A specific feature descriptor ($a\in A$)
$g(\cdot)$ Transformation mapping attributes to a subset of AA
$f_{\theta}$ Recommender system parameterized by $\theta$
Table 11. Impact of the backbone model used in AgenticTagger on downstream performance, Recall and NDCG.
Backbone LLM Sports Beauty CDs
R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10 R@5 N@5 R@10 N@10